I need help solving this:
x^2+xy+y^2=12
x+y=2

Answer by Alan3354(69443)

x^2+xy+y^2=12
x+y=2 --> y = 2-x
--------------------
Sub for y in the 1st eqn
x^2+xy+y^2=12
x^2 + x*(2-x) + (2-x)^2=12
Solve for x
etc